OpenStudy (anonymous):

Use identities (no calculators) to nd the exact value for (sin pi/4)(cos pi/2) + (cos pi/4)(sin pi/2)

OpenStudy (anonymous):

this ones the hardest of them all

OpenStudy (anonymous):

but if you have a look at it, see what expansion it looks like

OpenStudy (anonymous):

and it becomes pretty easy

OpenStudy (anonymous):

this = sin(pi/4 + pi)??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry pi/2 + pi/4

OpenStudy (anonymous):

=sin 3/4 pi = sin 75 degrees

OpenStudy (anonymous):

i/m lost now!!

OpenStudy (anonymous):

can sin 75 degrees be written in surd form?

OpenStudy (anonymous):

its not 75 :|

OpenStudy (anonymous):

but sin75 can be written in surd form

OpenStudy (anonymous):

no hold om 3/4 pi = 3/4 X 180 = 135 degrees ruyght?

OpenStudy (anonymous):

in which case sin 135 = 1 /sqrt 2

OpenStudy (anonymous):

you can use sum angles expansion to find exact value of sin75 sin75 = sin(30+45) = sin30cos45+cos30sin45

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my brains not working to well today !

OpenStudy (anonymous):

Mathematica returns zero after entering the problem expression. The left side of the + sign is zero because Sin[Pi]= zero. Same way for the right side.

OpenStudy (watchmath):

It is equal to \(\sin(\frac{\pi}{4}+\frac{\pi}{2})=\sin(3\pi/4)=\sqrt{2}/2\)

OpenStudy (anonymous):

(Sin pi/4) was processed as Sin(pi)/4 by Mathematica.

OpenStudy (anonymous):

Wolfram may be incorrect but they are reporting that \[\text{Sin}\left[\frac{3 \pi }{4}\right]=\frac{1}{\sqrt{2}}\]